OpenStudy (anonymous):

Which of the following is the equation of a line parallel to the line y = 4x + 1, passing through the point (5,1)? A. 4x + y = 19 B. 4x - y = 19 C. 4x + y = -19 D. -4x - y = 19

OpenStudy (anonymous):

b)

OpenStudy (anonymous):

two parallel has same slope....y=mx+c ..(m is slope) since our line is parallel to y=4x+1 hence our line has slope 4... so our line may be as y=4x+k; as our line pass through (5,1) hence it must satisfy our equation.. 1=4*5+k means k=-19 y=4x+k y=4x-19